# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 4.1. $L^{2}$-uniform bound of the logarithmic derivative $\mathsf{d}^{N}$ of $\mu^{N}$
Using (4.37) and noting $\mu_{R}=\mu\circ\pi_{\overline{S}_{R}}^{-1}$, we deduce that, for each $\mathscr{F}_{R}^{[1]}$-measurable $\mathsf{h}=\mathsf{f}\otimes\mathsf{g}\in\mathscr{D}_{\circ\mathrm{b}}^{[1]}$ such that $\mathsf{f}\in C_{0}^{\infty}(S_{R})$,
For a non-negative $f$ such that $\int_{\mathbb{R}^{d}}\mathsf{f}d{x}=1$, we consider the probability measure $\mathsf{f}\otimes 1\mu^{[1]}$. Then, from (4.42), we can easily show that $\mathsf{d}_{R}\mathsf{f}$ is an $\{\mathscr{F}_{R}^{[1]}\}$-martingale under $\mathsf{f}\otimes 1\mu^{[1]}$. Furthermore, $\mathsf{d}^{\mu}\in L_{\mathrm{loc}}^{2}(\mu^{[1]})$ from (4.35). Hence, (4.39) follows from the martingale convergence theorems. Equation (4.40) follows immediately from (4.39) and (4.42). Thus, $\mathsf{d}^{\mu}$ is the logarithmic derivative of $\mu$. From (4.40) and (4.42),
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}^{\mu}d\mu^{[1]}=\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}_{R}d\mu^{[1]}.$ (4.43)
Recall that $\mathsf{d}_{R}$ is $\mathscr{F}_{R}^{[1]}$-measurable. Then, from (4.43), we obtain (4.41). ∎
5. Explicit formulas for the logarithmic derivative $\mathsf{d}^{\mu}$: Proof of Theorem 2.1
In this section, we present explicit formulas for the logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$ and complete the proof of Theorem 2.1. Although Theorem 4.3 established the existence of a logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$, this result only confirms its existence and does not address the uniqueness of solutions to the associated ISDE or the existence of strong solutions.
This section assumes ((A1))((A3)). These assumptions are omitted from the statements of the Lemmas and Theorems for brevity.
The goal of this subsection is to construct the logarithmic derivative of $\mu_{R,\mathsf{y}}$, where $\mu_{R,\mathsf{y}}$ is the regular conditional probability defined by $\mu_{R,\mathsf{y}}=\mu(\cdot\mid\pi_{\overline{S}_{R}^{c}}(\mathsf{s})=\pi_{\overline{S}_{R}^{c}}(\mathsf{y}))$.
Let $\mu^{N}$ be the distribution of the $N$-particle system described in Section 1. For $\mu^{N}$-a.s. $\mathsf{y}\in\mathsf{S}$, let $\mu_{R,\mathsf{y}}^{N}$ be the regular conditional probability of $\mu^{N}$ given by
$\mu_{R,\mathsf{y}}^{N}=\mu^{N}(\cdot|\pi_{\overline{S}_{R}^{c}}(\mathsf{s})=\pi_{\overline{S}_{R}^{c}}(\mathsf{y})).$ (5.1)
be the one-reduced Campbell measure of $\mu_{R,\mathsf{y}}^{N}$. Then, from ((A1)) and ((A2)), $\rho_{R,\mathsf{y}}^{N,1}({x})$ is continuous in x{x} on $\overline{S}_{R}\cap O_{N}$. We denote by $\mathsf{d}_{R,\mathsf{y}}^{N}$ the logarithmic derivative of $\mu_{R,\mathsf{y}}^{N}$ for $\mu^{N}$-a.s. $\mathsf{y}$, where $\mathsf{d}_{R,\mathsf{y}}^{N}$ is defined by (4.10) by replacing $\mu_{R}^{N,[1]}$ by $\mu_{R,\mathsf{y}}^{N,[1]}$. Because $\rho_{R,\mathsf{y}}^{N,1}$ does not exist on $S_{R}^{c}$, $\mathsf{d}_{R,\mathsf{y}}^{N}$ is only defined on $\overline{S}_{R}\times\overline{\mathsf{S}}_{R}$, where $\overline{\mathsf{S}}_{R}$ is the configuration space over $\overline{S}_{R}$.
Let $\mathfrak{l}=(\mathfrak{l}^{i})$ be the label in (2.43). We set $\mathsf{y}^{N}=\mathsf{0}$ for $\mathsf{y}=\mathsf{0}$ and
$\mathsf{y}^{N}=\sum_{i=1}^{\mathsf{y}(\mathbb{R}^{d})\wedge N}\delta_{\mathfrak{l}^{i}(\mathsf{y})}\quad\text{ for $\mathsf{y}\neq\mathsf{0}$}.$ (5.3)
[…]

### v4

@@ 9.1. The $m$-labeled process for the upper Dirichlet form
From Lemma 7.7 ii, $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ is a quasi-regular, strongly local Dirichlet form on $L^{2}(\mu^{[m]})$. Hence, we obtain a $\mu^{[m]}$-symmetric diffusion $\overline{\mathbf{X}}^{[m]}$ properly associated with $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ on $L^{2}(\mu^{[m]})$ by the general theory of Dirichlet forms [5].
Because $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ is a quasi-regular Dirichlet form, we use the Fukushima decomposition (cf. [5, Th. 4.2.6]). We regard the coordinate function $x^{i}$ as a function on $(\mathbb{R}^{d})^{m}\times\mathsf{S}$ in an obvious manner for $1\leq i\leq m$. Then $x^{i}\in\mathscr{D}_{\mathrm{upr},\mathrm{loc}}^{[m]}$. Although $x^{i}\notin\mathscr{D}_{\mathrm{upr}}^{[m]}$, we can apply the Fukushima decomposition to $x^{i}$ by localization.
Consider the additive functional $A^{[x^{i}]}(t):=\overline{X}_{t}^{i}-\overline{X}_{0}^{i}$ of $\overline{\mathbf{X}}^{[m]}$. Let $\zeta$ be the lifetime of the diffusion $\overline{\mathbf{X}}^{[m]}$. Applying the Fukushima decomposition to $x^{i}$, we have a decomposition of $\overline{\mathbf{X}}^{[m]}$ such that, for $1\leq i\leq m$,
$\overline{X}_{t\wedge\zeta}^{i}-\overline{X}_{0}^{i}=M_{t\wedge\zeta}^{[x^{i}]}+N_{t\wedge\zeta}^{[x^{i}]},$ (9.6)
$M_{t\wedge\zeta}^{[x^{i}]}=\int_{0}^{t\wedge\zeta}\sigma(\overline{X}_{u}^{i})dB_{u}^{i}.$ (9.7)
For any $g\in\mathscr{D}_{\circ\mathrm{b}}^{[m]}$,
$-\mathscr{E}^{[m]}(x^{i},g)$ $=\int_{(\mathbb{R}^{d})^{m}\times\mathsf{S}}\mathfrak{b}(s^{i},\sum_{j\neq i}^{m}\delta_{s^{j}}+\sum_{k>m}\delta_{s^{k}})g(\mathsf{s}^{[m]})\mu^{[m]}(d\mathsf{s}^{[m]}).$ (9.8)
Here, we set $\mathsf{s}^{[m]}=(s^{1},\ldots,s^{m},\sum_{j>m}\delta_{s^{j}})$. Using (9.5) and (9.8) together with localization and [8, Th. 5.2.4], we have
$N_{t\wedge\zeta}^{[x^{i}]}=\int_{0}^{t\wedge\zeta}\mathfrak{b}(\overline{\mathscr{X}}_{u}^{i})du.$ (9.9)
Hence, from (9.6), (9.7), and (9.9), we obtain
$\overline{X}_{t\wedge\zeta}^{i}-\overline{X}_{0}^{i}=\int_{0}^{t\wedge\zeta}\sigma(\overline{X}_{u}^{i})dB_{u}^{i}+\int_{0}^{t\wedge\zeta}\mathfrak{b}(\overline{\mathscr{X}}_{u}^{i})du.$ (9.10)
Applying the Lyons–Zheng decomposition [5, p.284] to $x^{i}$, we see, by (9.10),
$\overline{X}_{t\wedge\zeta}^{i}-\overline{X}_{0}^{i}=\frac{1}{2}\Big\{\int_{0}^{t\wedge\zeta}\sigma(\overline{X}_{u}^{i})dB_{u}^{i}-\int_{0}^{t\wedge\zeta}\sigma(\overline{X}_{u}^{i})dB_{u}^{i}\circ r_{t}\Big\}.$ (9.11)
Let $\overline{\mathbf{X}}^{m}=(\overline{X}^{i})_{i=1}^{m}$ and $\overline{\mathsf{X}}_{t}^{m*}=\sum_{j>m}^{\infty}\delta_{\overline{X}_{t}^{j}}$. Then $\overline{\mathbf{X}}^{[m]}=(\overline{\mathbf{X}}^{m},\overline{\mathsf{X}}^{m*})$. Let
Then $\zeta=\zeta_{\mathrm{tp}}\wedge\zeta_{\mathrm{cnf}}$. Suppose $\zeta<\infty$. Taking $t\to\infty$ in (9.11) yields
$\overline{X}_{\zeta}^{i}-\overline{X}_{0}^{i}=\frac{1}{2}\Big\{\int_{0}^{\zeta}\sigma(\overline{X}_{u}^{i})dB_{u}^{i}-\int_{0}^{\zeta}\sigma(\overline{X}_{u}^{i})dB_{u}^{i}\circ r_{t}\Big\}.$
The right-hand side is finite. Hence, $\overline{X}_{\zeta}^{i}$ is finite. By $\lim_{t\to\zeta}\overline{X}_{t\wedge\zeta}^{i}=\overline{X}_{\zeta}^{i}$, $X^{i}$ does not explode at finite time $\zeta$. Hence, we obtain $\zeta=\zeta_{\mathrm{cnf}}$.
$\lim_{n\to\infty}\mathrm{Cap}^{\mu}(\mathsf{K}_{\mathsf{q}}^{c})=0.$ (9.12)
Let $\mathbf{a}=\{a_{\mathsf{q}}\}_{\mathsf{q}\in\mathbb{N}}$ and $\mathsf{K}[a_{\mathsf{q}}]$ be as in (13.3) and (13.4), respectively. By the property of compact sets of $\mathsf{S}$ [11], there exists such an $\{a_{\mathsf{q}}\}_{\mathsf{q}\in\mathbb{N}}$ satisfying
Hence, from (9.12) and (9.13),
$\mathrm{Cap}^{\mu}((\cup_{\mathsf{q}\in\mathbb{N}}\mathsf{K}[a_{\mathsf{q}}])^{c})=0.$ (9.14)
Using (9.14) and Lemma 9.1, we have
$\mathrm{Cap}^{[m]}((\mathfrak{u}^{[m,0]})^{-1}(\cup_{\mathsf{q}\in\mathbb{N}}\mathsf{K}[a_{\mathsf{q}}])^{c})=0.$
Hence, using the general theory of Dirichlet forms [8, 5], we obtain
$\overline{P}^{[m]}(\overline{\mathsf{X}}_{t}^{m*}\in(\mathfrak{u}^{[m,0]})^{-1}(\cup_{\mathsf{q}\in\mathbb{N}}\mathsf{K}[a_{\mathsf{q}}])^{c}\text{ for some }0\leq t<\infty)=0.$
Here $\overline{P}^{[m]}$ is the diffusion measure associated with $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ on $L^{2}(\mu^{[m]})$.
From this, we deduce $\zeta_{\mathrm{cnf}}=\infty$, $\overline{P}^{[m]}$-a.s. We have already proved that $\zeta=\zeta_{\mathrm{cnf}}$. Therefore, $\zeta=\infty$, $\overline{P}^{[m]}$-a.s. Thus $\overline{\mathbf{X}}^{[m]}$ is conservative, which implies i.
We obtain (9.1) from $\zeta_{\mathrm{tp}}=\zeta=\infty$. Since $\overline{\mathbf{X}}^{[m]}$ is $\mu^{[m]}$-symmetric and conservative by i, for all $R,T\in\mathbb{N}$, there exists a constant $c_{9.1}$ such that
$E\Big[\int_{0}^{T}\lvert\mathfrak{b}(\overline{\mathscr{X}}_{u}^{i})\rvert^{2}du;\overline{\mathbf{X}}_{0}^{m}\in S_{R}^{m}\Big]$ $\leq E\Big[\int_{0}^{T}\sum_{i=1}^{m}\lvert\mathfrak{b}(\overline{\mathscr{X}}_{u}^{i})\rvert^{2}du;\overline{\mathbf{X}}_{0}^{m}\in S_{R}^{m}\Big]$
[…]
From this and Proposition 6.3, we obtain (9.2). This proves ii.
By $\zeta=\infty$, iii follows from (9.10) and (9.11). This yields iii. ∎
Remark 9.1.
In the proof of Theorem 9.1, we applied the Lyons–Zheng decomposition directly to the limit process $\overline{\mathbf{X}}^{[m]}$, since the upper Dirichlet form $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ is quasi-regular by Lemma 7.7 ii.
At that point, however, the quasi-regularity of the lower Dirichlet form $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{lwr}}^{[m]})$ had not yet been established. For this reason, in the proof of Theorem 8.1 we instead applied the Lyons–Zheng decomposition to the finite-volume Dirichlet forms $(\mathscr{E}^{[m]}_{R},\underline{\mathscr{D}}_{R}^{[m]})$, which are known to be quasi-regular by Lemma 7.3.
9.2. Weak solutions of ISDEs for the upper Dirichlet form and the proof of Theorem 2.2
The proof in the present section is different from that of Subsection 8.2. The upper Dirichlet form Dirichlet form $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ on $L^{2}(\mu^{[m]})$ is quasi-regular. Hence, we will utilize the Dirichlet form theory and the general theorems developed in [26, 27].
For a label $\mathfrak{l}=(\mathfrak{l}^{i})_{i\in\mathbb{N}}$, let $\mathfrak{l}^{[0]}(\mathsf{s})=\mathsf{s}$ and $\mathfrak{l}^{[m]}(\mathsf{s})=((\mathfrak{l}^{i}(\mathsf{s}))_{i=1}^{m},\sum_{j>m}\delta_{\mathfrak{l}^{j}(\mathsf{s})})$. Let $\mathfrak{l}_{\mathrm{path}}\!:\!W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})\!\to\!W(\mathbb{R}^{d})^{\mathbb{N}}$ be as in Lemma 13.2. We define
Here, for $\mathbf{w}=(w^{i})$, we define the measure-valued path $\mathsf{w}$ by $\mathsf{w}_{t}=\sum_{i}\delta_{w_{t}^{i}}$. From Lemma 13.2, $\mathfrak{l}_{\mathrm{path}}(\mathsf{w})_{0}=\mathfrak{l}(\mathsf{w}_{0})$. Hence, $\mathfrak{l}_{\mathrm{path}}^{[m]}(\mathsf{w})_{0}=\mathfrak{l}^{[m]}(\mathsf{w}_{0})$.
Let $\overline{P}_{\mathfrak{l}^{[m]}(\mathsf{s})}^{[m]}$ be the diffusion measure of $\overline{\mathbf{X}}^{[m]}$ staring at $\mathfrak{l}^{[m]}(\mathsf{s})$ properly associated with $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ on $L^{2}(\mu^{[m]})$. By Lemma 7.7, $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ is a quasi-regular, strongly local symmetric Dirichlet form on $L^{2}(\mu^{[m]})$.
Lemma 9.2 ([26, Th. 2.4]).
For $\mu$-a.s. $\mathsf{s}$,
$\overline{P}_{\mathfrak{l}^{[m]}(\mathsf{s})}^{[m]}=\overline{P}^{\mu}_{\mathsf{s}}\circ(\mathfrak{l}_{\mathrm{path}}^{[m]})^{-1},\quad m\in\{0\}\cup\mathbb{N}.$ (9.15)
Proof.
Lemma 9.2 follows from the argument of [26, Th. 2.4]. We briefly explain below the difference between Lemma 9.2 and [26, Th. 2.4].
In [26, Th. 2.4], the following were assumed:
i $(\mathscr{E}^{[k]},\mathscr{D}_{\circ}^{[k]})$ is closable on $L^{2}(\mu^{[k]})$ for $k=0$ and $m$.
ii The $m$-point correlation function of $\mu$ is locally bounded for each $m\in\mathbb{N}$.
iii $\mathrm{Cap}(\mathsf{S}_{\mathrm{s}}^{c})=0$.
iv each tagged particle of $\overline{\mathsf{X}}$ does not explode.
In Lemma 9.2, i follows from Lemma 7.5.
In [26, Th. 2.4], ii was used only for proving the quasi-regularity of $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$. Thus, we can replace ii by Lemma 7.7.
In [26, Th. 2.4], iii was used to prove the non-collision of the tagged particles $\{\overline{X}^{i}\}_{i\in\mathbb{N}}$ of $\overline{\mathbf{X}}^{[m]}$ via [26, Lem. 4.1]. Hence, Theorem 9.1 ii applies.
iv also follows from Theorem 9.1 ii.
Thus, Lemma 9.2 can be proved by following the proof of [26, Th. 2.4]. ∎
For $\mathbf{w}=(w^{i})\in W(\mathbb{R}^{d})^{\mathbb{N}}$, let $\mathfrak{u}_{\mathrm{path}}^{[m]}(\mathbf{w})=(w^{1},\ldots,w^{m},\sum_{i>m}\delta_{w^{i}})$.
Theorem 9.2.
i There exists a family of probability measures $\overline{P}_{\mathbf{x}}^{\mathbb{N}}$ on $W(\mathbb{R}^{d})^{\mathbb{N}}$ such that, for $\mu$-a.s. $\mathsf{s}$,
$\quad\quad\quad\overline{P}_{\mathfrak{l}^{[m]}(\mathsf{s})}^{[m]}=\overline{P}_{\mathfrak{l}(\mathsf{s})}^{\mathbb{N}}\circ(\mathfrak{u}_{\mathrm{path}}^{[m]})^{-1},\quad\mathfrak{m}\in\{0\}\cup\mathbb{N}.$ (9.16)
ii Let $\overline{\mathbf{X}}=\mathbf{w}$. Then $\overline{\mathbf{X}}$ under $\overline{P}_{\varphi\mu\circ\mathfrak{l}^{-1}}^{\mathbb{N}}$ is a weak solution of (2.33) with initial distribution $\varphi\mu\circ\mathfrak{l}^{-1}$ that satisfies ((SIN)), ((NBJ)), and ((AC))μ.
Proof.
From (9.15), the family of distributions $\overline{P}_{\mathfrak{l}^{[m]}(\mathsf{s})}^{[m]}$ is consistent. Hence, using the Kolmogorov construction theorem, we obtain i.
Let $\overline{\mathbf{X}}^{[m]}$ be as in Theorem 9.1. Then its first $m$ components form a weak solution of (2.33). By the consistency (9.16), $\overline{\mathbf{X}}$ is a weak solution of (2.33).
From (9.3), for $\overline{\mathbf{X}}^{[m]}$, none of the $m$-labeled particles explode or collide with the other labeled particles for each $m\in\mathbb{N}$. Combining this with the consistency (9.16), each tagged particle $\overline{X}^{i}$ of $\overline{\mathbf{X}}=(\overline{X}^{i})_{i\in\mathbb{N}}$ neither explodes nor collides with the others. We have thus obtained ((SIN)).
We obtain ((NBJ)) from the martingale decomposition (9.4). The proof is analogous to that of Theorem 8.2 and is omitted.
Proof of Theorem 2.2. From Theorem 9.1, $\overline{\mathbf{X}}^{[m]}$ is a $\mu^{[m]}$-symmetric, conservative diffusion for each $m\in\{0\}\cup\mathbb{N}$. From Theorem 9.2, there exists a fully labeled process $\overline{\mathbf{X}}$ associated with $\overline{\mathbf{X}}^{[m]}$, and $\overline{\mathbf{X}}$ is a weak solution of (2.33) with initial distribution $\varphi\,\mu\circ\mathfrak{l}^{-1}$. Here we may take any function $\varphi$ satisfying (2.44). Using Fubini’s theorem, we obtain $\overline{\mathbf{X}}$ such that $\overline{\mathbf{X}}$ is a weak solution of (2.33) starting at $\mathfrak{l}(\mathsf{s})$ for $\mu$-a.s. $\mathsf{s}$. This completes the proof. ∎
10. Unique strong solutions of ISDEs: Proof of Theorem 2.3
To prove Theorem 2.3, we utilize a general result developed in [34, 18].
A key assumption for the uniqueness of solutions is the tail triviality of the associated RPFs. Starting from the tail decomposition of $\mu$, we introduce the tail decomposition of Dirichlet forms in Section 10.1. Using this, we complete the proof of Theorem 2.3 in Section 10.2.
Unless stated otherwise, all results in Section 7 are proved under Assumptions ((A1))((A3)).
10.1. Tail decomposition of Dirichlet forms and weak solution
Let $\mu_{\mathsf{a}}$, $\mathsf{a}\in\mathsf{S}_{0}$, be the tail decomposition of $\mu$ given in Definition 2.2. Then $\mu=\int\mu_{\mathsf{a}}d\mu(\mathsf{a})$ and $\mu(\mathsf{S}\backslash\mathsf{S}_{0})=0$. For all $R\in\mathbb{N}$, $\mu$-a.s. $\mathsf{a}$, and $\mu_{\mathsf{a}}$-a.s. $\mathsf{y}$, let
Let $\mu_{\mathsf{a},R,\mathsf{y}}^{[m]}$ be the $m$-reduced Campbell measure of $\mu_{\mathsf{a},R,\mathsf{y}}$. Let $\overline{\mathsf{S}}_{R}$ be the configuration space over $\overline{S}_{R}$ and $\overline{\mathsf{S}}_{R}^{k}=\{\mathsf{s}\in\overline{\mathsf{S}}_{R};\mathsf{s}(\overline{S}_{R})=k\}$.
Lemma 10.1.
For $\mu$-a.s. $\mathsf{a}$, $\mu_{\mathsf{a}}$-a.s. $\mathsf{y}$, and each $R\in\mathbb{N}$, the following hold:
i Let $m\in\{0\}\cup\mathbb{N}$. Then $\mu_{\mathsf{a},R,\mathsf{y}}^{[m]}$ has a density $\mathfrak{m}_{\mathsf{a},R,\mathsf{y}}^{[m]}$ with respect to $\Lambda_{R}^{[m]}$ such that $\mathfrak{m}_{\mathsf{a},R,\mathsf{y}}^{[m]}$ is bounded and continuous on $\overline{S}_{R}^{m}\times\overline{\mathsf{S}}_{R}^{k}$ for each $k\in\{0\}\cup\mathbb{N}$.
Proof.
Let $\mathsf{d}_{R,\mathsf{y}}$ and $\mathsf{d}_{\mathsf{a},R,\mathsf{y}}$ be the logarithmic derivatives of $\mu_{R,\mathsf{y}}$ and $\mu_{\mathsf{a},R,\mathsf{y}}$, respectively. From Proposition 5.3, $\mathsf{d}_{R,\mathsf{y}}({x},\mathsf{s})=\mathsf{d}_{\mathsf{a},R,\mathsf{y}}({x},\mathsf{s})$ for $\mu_{\mathsf{a}}^{[1]}$-a.e. $({x},\mathsf{s})$ for $\mu$-a.s. $\mathsf{a}$. Hence, from Theorem 5.1 i,
$\mathsf{d}_{\mathsf{a},R,\mathsf{y}}({x},\mathsf{s})=\beta\Big(-\nabla\Phi({x})-\sum_{s^{i}\in\overline{S}_{R}}\nabla\Psi({{x}-s^{i}})+\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i}}{x}^{\mathbf{i}}+\mathscr{R}_{R,\mathsf{y}}^{\ell}({x})\Big).$
In Lemma 7.1, the existence of a bounded continuous density $\mathfrak{m}_{R,\mathsf{y}}^{[m]}$ of $\mu_{R,\mathsf{y}}^{[m]}$ was derived from the explicit representation (5.30) of $\mathsf{d}_{R,\mathsf{y}}$. Using the above representation of $\mathsf{d}_{\mathsf{a},R,\mathsf{y}}$, assertion i follows by the same argument.
Assertion ii can be proved analogously to Lemma 7.1 ii. ∎
Lemmas 10.210.7 below correspond to Lemmas 7.27.7, respectively. The proof of Lemma 10.kk is similar to that of Lemma 7.kk for $1\leq k\leq 6$; therefore, we omit those proofs.
Let $\mu_{\mathsf{a}}^{[m]}$ be the $m$-reduced Campbell measure of $\mu_{\mathsf{a}}$. By replacing $\mu^{[m]}$ with $\mu_{\mathsf{a}}^{[m]}$ and $\mu_{\mathsf{a},R,\mathsf{y}}^{[m]}$ in (7.2), we define $(\mathscr{E}_{\mathsf{a},R}^{[m]},\mathscr{D}_{\mathsf{a},R\bullet}^{[m]})$ and $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\mathscr{D}_{\mathsf{a},R,\mathsf{y}\bullet}^{[m]})$, respectively.
Lemma 10.2.
For $\mu$-a.s. $\mathsf{a}$, $R\in\mathbb{N}$, and $m\in\{0\}\cup\mathbb{N}$, the following hold:
i $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\mathscr{D}_{\mathsf{a},R,\mathsf{y}\bullet}^{[m]})$ is closable on $L^{2}(\mu_{\mathsf{a},R,\mathsf{y}}^{[m]})$ for $\mu_{\mathsf{a}}$-a.s. $\mathsf{y}$.
$i$ $(\mathscr{E}_{\mathsf{a},R}^{[m]},\mathscr{D}_{\mathsf{a},R\bullet}^{[m]})$ is closable on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
iii $(\mathscr{E}_{\mathsf{a},R}^{[m]},\mathscr{D}_{\mathsf{a},R\circ}^{[m]})$ is closable on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
Let $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ be the closure of $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\mathscr{D}_{\mathsf{a},R,\mathsf{y}\bullet}^{[m]})$ on $L^{2}(\mu_{\mathsf{a},R,\mathsf{y}}^{[m]})$. Let $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$ and $(\mathscr{E}_{\mathsf{a},R}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$ be the closures of $(\mathscr{E}_{\mathsf{a},R}^{[m]},\mathscr{D}_{\mathsf{a},R\bullet}^{[m]})$ and $(\mathscr{E}_{\mathsf{a},R}^{[m]},\mathscr{D}_{\mathsf{a},R\circ}^{[m]})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$, respectively.
Lemma 10.3.
For $\mu$-a.s. $\mathsf{a}$, $R\in\mathbb{N}$, and $m\in\{0\}\cup\mathbb{N}$, the following hold:
i $(\mathscr{E}_{\mathsf{a},R,\mathsf{y}}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R,\mathsf{y}}^{[m]})$ is a strongly local, quasi-regular Dirichlet form on $L^{2}(\mu_{\mathsf{a},R,\mathsf{y}}^{[m]})$ for $\mu$-a.s. $\mathsf{y}$.
$i$ $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$ is a strongly local, quasi-regular Dirichlet form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
iii $(\mathscr{E}_{\mathsf{a},R}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$ is a strongly local closed form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
Let $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$ be the increasing limit of $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$, $R\in\mathbb{N}$.
Lemma 10.4.
For $\mu$-a.s. $\mathsf{a}$ and $m\in\{0\}\cup\mathbb{N}$, the following hold:
i The form $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$ is a closed form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$ and is the strong resolvent limit of $(\mathscr{E}_{\mathsf{a},R}^{[m]},\underline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
$i$ $(\mathscr{E}_{\mathsf{a}}^{[m]},\cup_{R\in\mathbb{N}}\overline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$ is closable on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. The closure $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ is the limit of $(\mathscr{E}_{\mathsf{a}}^{[m]},\overline{\mathscr{D}}_{\mathsf{a},R}^{[m]})$ in the strong resolvent sense on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
Let $\mathscr{D}_{\mathsf{a}\bullet}^{[m]}=\{f\in C_{0}^{\infty}((\mathbb{R}^{d})^{m})\otimes\mathscr{D}_{\bullet};\mathscr{E}_{\mathsf{a}}^{[m]}(f,f)<\infty,f\in L^{2}(\mu_{\mathsf{a}}^{[m]})\}$. We define $\mathscr{D}_{\mathsf{a}\circ}^{[m]}$ by replacing $\mathscr{D}_{\bullet}$ with $\mathscr{D}_{\circ}$.
Lemma 10.5.
For $\mu$-a.s. $\mathsf{a}$ and $m\in\{0\}\cup\mathbb{N}$, the following hold:
i $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a}\bullet}^{[m]})$ is closable on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
$i$ $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a}\circ}^{[m]})$ is closable on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
Let $(\mathscr{E}_{\mathsf{a}}^{[m]},\overline{\mathscr{D}_{\mathsf{a}\bullet}^{[m]}})$ and $(\mathscr{E}_{\mathsf{a}}^{[m]},\overline{\mathscr{D}_{\mathsf{a}\circ}^{[m]}})$ be the closure of $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a}\bullet}^{[m]})$ and $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a}\circ}^{[m]})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$, respectively.
Lemma 10.6.
For $\mu$-a.s. $\mathsf{a}$ and each $m\in\{0\}\cup\mathbb{N}$, the following hold:
i $(\mathscr{E}_{\mathsf{a}}^{[m]},\overline{\mathscr{D}_{\mathsf{a}\bullet}^{[m]}})=(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$.
$i$ $(\mathscr{E}_{\mathsf{a}}^{[m]},\overline{\mathscr{D}_{\mathsf{a}\circ}^{[m]}})=(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$.
Lemma 10.7.
For $\mu$-a.s. $\mathsf{a}$ and $m\in\{0\}\cup\mathbb{N}$, the following hold:
i $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$ is a strongly local Dirichlet form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
$i$ $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ is a strongly local, quasi-regular Dirichlet form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
Similarly as $\mathsf{d}_{R,\mathsf{y}}$ for $\mu_{R,\mathsf{y}}^{[1]}$ in Lemma 5.6, we obtain $\mathsf{d}_{\mathsf{a},R,\mathsf{y}}$ for $\mu_{\mathsf{a},R,\mathsf{y}}^{[1]}$. In Lemma 10.8, we prove that $\mathsf{d}_{\mathsf{a},R,\mathsf{y}}$ coincides with $\mathsf{d}^{\mu}$ under $\mu_{\mathsf{a},R,\mathsf{y}}^{[1]}$.
Lemma 10.8.
For $\mu$-a.s. $\mathsf{a}$, all $R\in\mathbb{N}$, and $\mu_{\mathsf{a}}$-a.s. $\mathsf{y}$, $\mathsf{d}_{\mathsf{a},R,\mathsf{y}}$ satisfies
Proof.
By $\mu_{\mathsf{a},R,\mathsf{y}}^{[1]}\mu_{\mathsf{a}}(d\mathsf{y})\mu(d\mathsf{a})=\mu^{[1]}$, we obtain, for $h\in C_{0}^{\infty}(S_{R})\otimes\mathscr{D}_{\circ\mathrm{b}}$,
This implies (10.1). ∎
[…]